# TaskCLIP: Extend Large Vision-Language Model for Task Oriented Object Detection

### v1

@@ Abstract
Task-oriented object detection aims to find objects suitable for accomplishing specific tasks. As a challenging task, it requires simultaneous visual data processing and reasoning under ambiguous semantics. Recent solutions are mainly all-in-one models. However, the object detection backbones are pre-trained without text supervision. Thus, to incorporate task requirements, their intricate models undergo extensive learning on a highly imbalanced and scarce dataset, resulting in capped performance, laborious training, and poor generalizability. In contrast, we propose TaskCLIP, a more natural two-stage design composed of general object detection and task-guided object selection. Particularly for the latter, we resort to the recently successful large Vision-Language Models (VLMs) as our backbone, which provides rich semantic knowledge and a uniform embedding space for images and texts. Nevertheless, the naive application of VLMs leads to sub-optimal quality, due to the misalignment between embeddings of object images and their visual attributes, which are mainly adjective phrases. To this end, we design a transformer-based aligner after the pre-trained VLMs to re-calibrate both embeddings. Finally, we employ a trainable score function to post-process the VLM matching results for object selection. Experimental results demonstrate that our TaskCLIP outperforms the state-of-the-art DETR-based model TOIST by 3.5% and only requires a single NVIDIA RTX 4090 for both training and inference.
@@ 1 Introduction
Objection detection algorithms have seen tremendous progress on datasets like COCO [23] and Pascal VOC [9], where they identify object instances of predefined categories in a scene [12]. However, in real-world applications, artificial intelligence is expected to handle a more specific “task-oriented object detection”. To this end, researchers propose a new benchmark named COCO-Tasks [33] that is much more challenging than traditional objection detection tasks. Fig. 1(a) highlights that a simple, yet practical task never confines itself to a closed and predefined set of objects (or text prompts consisting of nouns). The inputs are in the form of affordance (e.g., “Sit Comfortably”), conditioning the selection of objects on their capability to fulfill a task. Notably, this brings a major challenge to prior object detection algorithms. Affordance as the prompt is intrinsically ambiguous, e.g., best suitable objects range from “Bed” to the seemingly unrelated “Toilet” for the same task “Sit Comfortably”, depending on different scenes.
Current solutions follow either a two-stage or single-stage design to tackle this challenge. The former starts with regular object detection, followed by task-driven object selection, whereas the latter aims to achieve both sub-tasks with one single model. The two-stage design proposed in [33] applies object preference ranking on top of objects extracted by traditional detectors. More recent one-stage designs [20] mainly resort to advanced object detection models like the transformer-based DETR [3] as the backbone. Work [36] further leverages Large Language Models (LLMs) to generate intermediate visual affordance features, which close the semantic gap between affordances that are mainly verb phrases and object types that are nouns.
However, as intriguing as the all-in-one model sounds, the costs and difficulties behind it cannot be ignored. As pointed out by the prior solution [3], transformers are considered data hungry while obtaining datasets like COCO-Tasks remains extremely difficult. As shown in Fig. 1(b), the dataset also suffers from severe class imbalance, since only a handful of objects are relevant to the tasks concerned and the rest are labeled as negative samples. In addition, all prior works require an explicit input of task-specific information to the object detection or selection model , fundamentally limiting their robustness against dataset domain shift or even slight changes in the affordance prompts. This leads to our question on the optimality of training these task-specialized one-stage models from scratch, given the data scarcity in this field.
@@
Our pursuit of solutions takes cues from the reasoning process of human beings, which is naturally multi-stage. Initially, they identify potential instances of objects and extract their visual features. Simultaneously, humans parse the task, discerning the required visual or functional attributes of candidate objects that are essential for task completion. Subsequently, they evaluate instances against these task-relevant attributes, selecting instances that align most closely with the task’s requirements. Crucially, humans do not merely seek instances directly matching the task; instead, they engage in a step-by-step reasoning process that leverages the common features of suitable objects as clues. This is similar in spirit to existing two-stage solutions. Yet their performance is suboptimal mainly due to the bottleneck in the second stage, i.e., the association of object visual features and the semantics in task affordance.
@@
In this paper, we present TaskCLIP, a two-stage task-oriented object detection algorithm that tackles challenges in prior algorithms all at once, motivated by the recent success of large-scale Vision-Language Models (VLMs) on zero-shot transfer for language-supervised vision tasks. In the first stage, we perform general object detection for each scene. Meanwhile, we parse the task affordance, via LLMs[2, 29], into several adjectives that describe common attributes of suitable objects. In our second stage, VLMs like CLIP [31] and Flamingo [1] serve as the pre-trained backbones, which match image and text embeddings. The resulting affinity matrices then guide the selection of suitable objects. They enable the alignment of visual and text features by constructing a unified and high-quality embedding space, thereby resolving the bottleneck in current two-stage designs. More importantly, we avoid the complex customization, costly training, and poor generalizability of all-in-one models, where all parameters are learned from scratch and tuned specifically for a few tasks in COCO-Tasks with limited training samples.
@@
Nevertheless, incorporating VLMs like CLIP into our two-stage design is non-trivial. Standard multimodal VLMs, while proficient in alignment, fall short of the reasoning capabilities required for selecting the most suitable objects. On one hand, these models primarily focus on aligning nouns with visual embeddings, while abstract concepts, such as phrases with adjectives, are poorly supported (as shown in Fig. 1(c)). On the other hand, large VLMs are trained on rich and diverse natural language annotations that contain not only nouns for objects. In response, we propose a re-alignment mechanism that blends seamlessly into the end-to-end model training. It re-enables the multimodal VLM to encode adjectives in the common embedding space and properly align them with corresponding visual embeddings.
@@
  • Instead of training multi-modal model from scratch, we propose the first VLM-based, two-stage, task-oriented object detection framework, leveraging the extensive semantic information from vision-language pretraining and its capability to support a calibrated joint vision-text embedding space.
@@
  • Our framework provides natural task-driven reasoning and high-quality object detection, while also demonstrating higher training efficiency than previous works. Compared to the DETR-based models, our TaskCLIP achieves up to 3.5% improvement in mAP@0.5 on the COCO-Tasks dataset, requiring only a single RTX 4090 GPU to support both training and inference.
@@ 2.1 Task-oriented Object Detection
This focuses on identifying the most suitable objects within an image to accomplish a specific task, such as serving wine. It requires the ability to reason and select objects based on a deep understanding of the task and the visual scene. Thus, task-oriented object detection is much more complex than simple visual object detection [8, 7, 32, 24] and scene understanding [6, 39, 34, 5]. The first task-driven object detection dataset, named COCO-Tasks, is proposed by a study [33] based on MS COCO 2014 [23]. It also introduced a two-stage framework as the benchmark, which initially employs an existing object detection network to identify potentially suitable items, followed by gated graph neural networks (GGNN) [21] to model global inter-object relation for object preference ranking. Subsequent works, TOIST and CoTDet [20, 36], are based on DETR [42, 19, 3] and can be considered single-stage frameworks. Compared to GGNN, the self-attention module in DETR-based models is more effective in terms of integrating text and visual information. However, training DETR-based models for task-oriented object detection faces multiple challenges due to the complex model structure and a large number of parameters. On the contrary, our TaskCLIP resorts to vision-language pertaining and offers a more efficient alternative.
@@ 2.2 Vision Language Model
Extending pre-trained VLMs to support new applications has demonstrated their success in recent years. Among the most successful efforts, Flamingo [1], OpenAI CLIP [31], and OpenCLIP [4] have exhibited impressive performance in handling image-text matching, owing to their profound semantic knowledge and comprehensive understanding of content that spans both modalities. These models have been applied successfully in downstream applications such as object detection [35], action recognition [37], image captioning [28], dense prediction [41], and semantic segmentation [22]. However, existing CLIP-based algorithms primarily focus on matching image patches with nouns in text. Thus, additional calibration is necessary to facilitate the matching between image patches and visual attributes that are adjective phrases.
@@ 3 Method
Figure 2: TaskCLIP model architecture: (a) The overall framework of the design. (b) The architecture of the aligner module. (c) Detailed design of score function.
@@ 3.1 Task visual attributes preparation
Figure 3: TaskCLIP LLM prompt.
To close the semantic and reasoning gap between task affordance and suitable object class, we utilize an LLM to produce task-relevant visual common attributes, as is shown in Fig. 2(a)
1
1
[…]
. In Fig. 3, we delineate the prompts employed to direct the LLM in generating the appropriate visual attributes for each task. To illustrate this process, we consider task 10 (Serve wine) as an example. For task 10, the procedure unfolds as follows:
@@
(2). Building upon this initial response, we proceed to the second prompt: “Summarize the visual features of the [wine glass, cup, stemware, …] for serve wine.” It is designed to guide the LLM in summarizing the visual attributes of the mentioned items specifically in the context of the “serve wine” task.
(3). Ultimately, we obtain the common visual attributes of objects that can afford this task from LLM responses, such as “narrow rim” and “transparency”. In total, we generate $\mathbf{N_{word}}=10$ common attributes.
@@ 3.2 Text and visual embedding vector generation
Meanwhile, we generate $N_{bbx}$ bounding boxes for all objects in the scene as is shown in Fig. 2(a)
2
2
[…]
. During the training process, we directly use ground truth bounding boxes, whereas, for inference, we use a pre-trained object detection network (such as Faster R-CNN [11] or YOLOv8 [18]). Notice that, in the basic setting of our framework during inference, the object detection network is only responsible for generating bounding boxes for all items in the image. We will introduce a grouping mechanism in section 3.5 to further leverage object classification information and guide our model’s final decision. Based on the bounding box coordination, we extract $N_{bbox}$ image patches. After generating $N_{word}$ common visual attributes and $N_{bbox}$ image patches, we pass them into pre-trained VLMs (such as OpenAI CLIP [31]) to generate text and image embeddings as is shown in Fig. 2(a)
3
3
[…]
. Here we use $L_{word}$ and $L_{bbox}$ to represent the lists of visual feature words and image patches. The lengths of the lists are $N_{word}$ and $N_{bbox}$, respectively. The computation process can be summarized as follows:
@@
Suppose the embedding dimension is D, the shape of the generated vision embedding matrix and text embedding matrix gonna be $\textbf{N}_{bbox}$$\times$×D and $\textbf{N}_{word}$$\times$×D.
@@ 3.3 Vision and text embedding space recalibration
Although OpenAI CLIP and its derivative works [31, 40] achieve high accuracy in matching text with image patch, they focus on noun (such as “wine glass” and “cup”) instead of visual attributes (such as “transparency” and “narrow rim”). To match the image patches with task-specific visual attribute words, it is pivotal to recalibrate the embedding space of vision and text. Therefore we design a transformer-based aligner module, next to the pre-trained VLM. In Fig. 2(a)
4
4
[…]
, we pass the vision embedding matrix and text embedding matrix into a multi-layer aligner module to generate the new recalibrated vision and text embedding matrix, $\textbf{E}^{\textbf{m}}_{\textbf{V}}$ and $\textbf{E}^{\textbf{m}}_{\textbf{T}}$ respectively. Here m represents the number of layers in aligner module. In Fig. 2(b), we present the network architecture of each layer of the aligner module. There are multiple layers of aligner module where each layer has two self-attention modules, one cross-attention module, and one feed-forward network layer. For the $\textbf{i}^{\textbf{th}}$ aligner module layer, the input ($\textbf{E}^{\textbf{i}}_{\textbf{V}}$ and $\textbf{E}^{\textbf{i}}_{\textbf{T}}$) gonna be the vision and text embedding matrix coming from last aligner module layer. At each layer, both vision and text embedding matrix will first go through a self-attention layer. The purpose of the self-attention layer is to learn the most optimal affordance that could be used to solve this task. For example, as is shown in Fig. 1(a) Scene 5, to solve the task of sit comfortably we prefer to choose chair over bed. After self-attention layer, we use a cross-attention layer to recalibrate the text embedding and vision embedding. The last part of the aligner network is a fully connected feed forward layer where the text embedding matrix will pass through. Like previous works that try to fine-tune pre-trained vision-language model [10, 14], in our aligner module, we focus on the text embedding vectors transformation. Therefore, we only pass text embedding matrix through the last feed forward layer.

### v2

@@ Abstract
Task-oriented object detection aims to find suitable objects for performing specific tasks. As a challenging task, it requires simultaneous visual data processing and reasoning under ambiguous semantics. Recent solutions are mainly all-in-one models. However, the object detection backbones are pre-trained without text supervision. Thus, to incorporate task requirements, their intricate models undergo extensive learning on a highly imbalanced and scarce dataset, resulting in capped performance, laborious training, and poor generalizability. In contrast, we propose TaskCLIP, a more natural two-stage design composed of general object detection and task-reasoning object selection. Particularly for the latter, we resort to the recently successful large Vision-Language Models (VLMs) as our backbone, which provides rich semantic knowledge and a uniform embedding space for images and texts. Nevertheless, the naive application of VLMs leads to suboptimal quality, due to the misalignment between embeddings of object images and their visual attributes, which are mainly adjective phrases. To this end, we design a transformer-based aligner after the pre-trained VLMs to recalibrate both embeddings. Finally, we employ a trainable score function to post-process the VLM matching results for object selection. Experimental results demonstrate that our TaskCLIP outperforms the DETR-based model TOIST in both accuracy ($+6.2\%$) and efficiency.
@@ 1 Introduction
Object detection algorithms have seen tremendous progress on datasets like COCO [29] and Pascal VOC [9], where they identify object instances of predefined categories in a scene [13, 17, 49]. However, in real-world applications, artificial intelligence is expected to handle more specific “task-oriented object detection”. To this end, researchers propose a new benchmark named COCO-Tasks [41] that is much more challenging than traditional object detection tasks. Fig. 1(a) highlights that a simple, yet practical task never confines itself to a closed and predefined set of objects (or text prompts consisting of nouns). The inputs are in the form of affordance (e.g., “Sit Comfortably”), conditioning the selection of objects on their capability to fulfill a task. In particular, this presents a major challenge for prior object detection algorithms. The affordability as a prompt is intrinsically ambiguous, for example, the best suitable objects range from “Bed” to the seemingly unrelated “Toilet” for the same task “Sit Comfortably”, depending on different scenes.
Current solutions follow either a two-stage or a single-stage design to tackle this challenge. The former starts with regular object detection, followed by task-driven object selection, whereas the latter aims to achieve both sub-tasks with one single model. The two-stage design proposed in [41, 16] applies a ranking of objects preference on top of objects extracted by traditional detectors. More recent one-stage designs [26] mainly resort to advanced object detection models such as the transformer-based DETR [3] as the backbone. Work [44] further leverages Large Language Models (LLMs) to generate intermediate visual affordance features, which close the semantic gap between affordances that are mainly verb phrases and object types that are nouns.
However, as intriguing as the all-in-one model sounds, the costs and difficulties behind it cannot be ignored. As pointed out by the previous solution [3], transformers are considered data hungry, while obtaining data sets such as COCO-Tasks remains extremely difficult. As shown in Fig. 1(b), the dataset also suffers from severe class imbalance, since only a handful of objects are relevant to the tasks concerned and the rest are labeled as negative samples. In addition, all prior works require an explicit input of task-specific information to the object detection or selection model [41, 26], fundamentally limiting their robustness against dataset domain shift or even slight changes in the affordance prompts. For instance, altering the task name to synonymous terms will cause a notable decline in the model’s performance, even if the underlying characteristics of the task remain unchanged. This brings us to the question of whether it is optimal to train these specialized one-stage models for specific tasks from scratch, considering the limited data available in this domain. It also raises another important issue: how can we design more versatile computer vision models capable of addressing a variety of tasks?
@@
Our pursuit of solutions takes cues from the reasoning process of human beings, which is naturally multi-stage. Initially, they identify potential instances of objects and extract their visual features. During the process, humans parse the task, discerning the visual or functional attributes required of the candidate objects that are essential for task completion. Subsequently, they evaluated instances against these tasks-relevant attributes, selecting instances that align the most closely with the task requirements. Crucially, humans do not simply seek instances directly matching the task; instead, they engage in a step-by-step reasoning process that takes advantage of the common features of suitable objects as clues. This is similar in spirit to existing two-stage solutions. However, their performance is suboptimal mainly due to the bottleneck in the second stage, i.e., the association of object visual features and the semantics in task affordance.
@@
In this paper, we present TaskCLIP, a two-stage task-oriented object detection algorithm that tackles challenges in prior algorithms all at once, motivated by the recent success of large-scale Vision-Language Models (VLMs) on zero-shot transfer for language-supervised vision tasks. In the first stage, we perform general object detection [12, 39] for each scene. Meanwhile, we parse the task affordance, via LLMs[2, 36], into several adjectives that describe common attributes of suitable objects. In our second stage, VLMs like CLIP [38] and Flamingo [1] serve as pre-trained backbones, which match image and text embeddings. The resulting affinity matrices then guide the selection of suitable objects. They enable the alignment of visual and text features by constructing a unified and high-quality embedding space, thereby resolving the bottleneck in current two-stage designs. More importantly, we avoid the complex customization, costly training, and poor generalizability of all-in-one models, where all parameters are learned from scratch and tuned specifically for a few tasks in COCO-Tasks with limited training samples. Given that TaskCLIP addresses the task through the selection of items meeting visual characteristics, our model can be utilized for other tasks not included in the training group, provided they share the same fundamental attributes.
@@
However, incorporating VLMs like CLIP into our two-stage design is not trivial. Standard multimodal VLMs, while proficient in alignment, fall short of the reasoning capabilities required to select the most suitable objects. On one hand, these models primarily focus on aligning nouns with visual embeddings, while abstract concepts, such as phrases with adjectives, are poorly supported (as shown in Fig. 1(c)). On the other hand, large VLMs are trained on rich and diverse natural language annotations that contain not only nouns for objects. In response, we propose a re-alignment mechanism that blends seamlessly into the end-to-end model training. It reinstates the multimodal VLM’s ability to encode adjectives within the shared embedding space and accurately associate them with the related visual embeddings.
@@
  • Instead of training a multi-modal model from scratch, we propose the first VLM-based, two-stage, task-oriented object detection framework, leveraging the extensive semantic information from vision-language pretraining and its capability to support a calibrated joint vision-text embedding space.
@@
  • Our framework provides natural task-driven reasoning and high-quality object detection, while also demonstrating higher training efficiency than previous works. In contrast to earlier DETR-based models, TaskCLIP attains similar accuracy on the COCO-Tasks dataset, yet exhibits significantly improved training efficiency and better generalizability.
@@ 2 Related Work
Figure 2: (a) CoTDet [44] workflow. (b). TaskCLIP (this work) workflow.
@@ 2.1 Task-oriented Object Detection
This focuses on identifying the most suitable objects within an image to accomplish a specific task, such as serving wine. It requires the ability to reason and select objects based on a deep understanding of the task and the visual scene. Thus, task-oriented object detection is much more complex than simple visual object detection [8, 7, 40, 30] and scene understanding [6, 48, 42, 5, 20, 21]. The first task-driven object detection dataset, named COCO-Tasks, is proposed by a study [41] based on MS COCO 2014 [29]. It also introduced a two-stage framework as a benchmark, which initially employs an existing object detection network to identify potentially suitable items, followed by gated graph neural networks (GGNN) [27] to model global inter-object relationship for object preference ranking. Subsequent works, TOIST and CoTDet [26, 44], are based on DETR [52, 24, 3] and can be considered as single-stage frameworks. Compared to GGNN, the self-attention module in DETR-based models is more effective in terms of integrating text and visual information. However, training DETR-based models for task-oriented object detection faces multiple challenges due to the complex model structure and the large number of parameters. In contrast, our TaskCLIP resorts to the pertaining vision language and offers a more efficient alternative. A further problem is that while previous DETR-based models include text modalities, they do not take into account the model’s ability to generalize. Consider CoTDet [44], as shown in Figure 2.(a). For each task, after generating task-specific visual affordance and visual feature embeddings using the LLM (ChatGPT) and the text encoder (RoBERTa [31]), the model requires fine-tuning of these embedding vectors. A change in the task name results in a significant decrease in detection accuracy.
@@ 2.2 Vision Language Model
Extending pre-trained VLMs to support new applications has demonstrated their success in recent years. Among the most successful efforts, Flamingo [1], OpenAI CLIP [38], and OpenCLIP [4] have exhibited impressive performance in handling image-text matching due to their deep semantic knowledge and a comprehensive understanding of content that spans both modalities. These models have been applied successfully in downstream applications such as object detection [43], action recognition [45], image captioning [35], dense prediction [51], disease prognosis [34], and semantic segmentation [28, 46, 25]. However, existing CLIP-based algorithms primarily focus on matching image patches with nouns in text. Thus, additional calibration is necessary to facilitate the matching between image patches and visual attributes that are adjective phrases.
@@ 3 Method
Figure 3: TaskCLIP model architecture: (a) The overall framework of the design. (b) The architecture of the aligner module. (c) Detailed design of the score function
@@ 3.1 Task visual attributes preparation
Table 1: TaskCLIP LLM prompt.
(a) Description words bank preparation
User: Please list items to solve the task of <Task Name>?
LLM response: [item1, item2, …]
User: Summarize visual featuresof the [item1,item2,…] for <Task Name>.
LLM response: [feature1, feature2, …]
(b) Synonym task generation
User: Please generate a synonym for <Task Name>
LLM response: <Task Name Synonym>
(c) Visual features selection
User: Please select 20 visual feature words from words bank for <Task Name>
LLM response: [feature1, feature2, ….]
To close the semantic and reasoning gap between task affordance and suitable object class, we utilize an LLM to produce task-relevant visual common attributes, as is shown in Fig. 2(b). In Table 1.(a), we delineate the prompts used to guide the LLM in generating the appropriate visual attributes for each task. To illustrate this process, we consider task 10 (Serve wine) as an example. For task 10, the procedure unfolds as follows:
@@
(2). Building on this initial response, we proceed to the second prompt: “Summarize the visual features of the [wine glass, cup, stemware, …] for serve wine.” It is designed to guide the LLM in summarizing the visual characteristics of the mentioned items specifically in the context of the “serve wine” task.
(3). Ultimately, we obtain the common visual attributes of objects that can afford this task from LLM responses, such as “narrow rim” and “transparency”. In total, we generate $\mathbf{N_{word}}=20$ common attributes.
(4). We save these generated common attributes in the description words bank. As depicted in Figure 2, when new tasks arise, the LLM will pick the most suitable attribute from the description words bank. Since TaskCLIP is trained by aligning item images with visual feature description words, even if the task name changes, as long as the intrinsic attributes of the new task can be articulated using the words from the bank, our model remains effective. Hence, TaskCLIP exhibits superior task generalizability compared to prior works.
@@ 3.2 Text and visual embedding vector generation
Meanwhile, we generate $N_{bbx}$ bounding boxes for all objects in the scene as is shown in Fig. 3(a). During the training process, we directly use ground truth bounding boxes, whereas, for inference, we use a pre-trained object detection network (such as Faster R-CNN [12] or YOLOv8 [23]). Notice that in the basic setting of our framework during inference, the object detection network is only responsible for generating bounding boxes for all items in the image. We will introduce a grouping mechanism in Section 3.6 to further leverage object classification information and guide our model’s final decision. Based on the coordination of the bounding box, we extract $N_{bbox}$ image patches. After generating $N_{word}$ common visual attributes and $N_{bbox}$ image patches, we pass them into pre-trained VLMs (such as OpenAI CLIP [38]) to generate text and image embeddings as is shown in Fig. 3(a). Here we use $L_{word}$ and $L_{bbox}$ to represent the lists of visual feature words and image patches. The lengths of the lists are $N_{word}$ and $N_{bbox}$, respectively. The computation process is summarized as:
@@
Suppose the embedding dimension is D, the shape of the generated vision embedding matrix and text embedding matrix is $\textbf{N}_{bbox}$$\times$×D and $\textbf{N}_{word}$$\times$×D. To better introduce the new knowledge into the vision language models, we also add vision and text adapter [11] after the CLIP model.
$\mathbf{E^{0}_{V}}=(1-\alpha)\times\mathbf{E^{0}_{V}}+ReLU(\mathbf{E^{0T}_{V}}\mathbf{W_{1}^{v}})\mathbf{W_{2}^{v}}$ (3)
$\mathbf{E^{0}_{T}}=(1-\beta)\times\mathbf{E^{0}_{T}}+ReLU(\mathbf{E^{0T}_{T}}\mathbf{W_{1}^{t}})\mathbf{W_{2}^{t}}$ (4)
Here $W_{1}^{v}$, $W_{2}^{v}$ and $W_{1}^{t}$, $W_{2}^{t}$ are the weights of bottleneck linear layers for vision and text adapter networks. $\alpha$ and $\beta$ are hyper-parameters, here we choose 0.3.
3.3 Global attention
To gain a clearer understanding of the role of each bounding box item within the overall scene, we initially encode the input image using the CLIP vision encoder. Subsequently, we use the embedding of the bounding box list ($E^{0}_{V}$) as the target and the image embedding as the memory. After the cross-attention module, we get the bounding box embedding with global information.
@@ 3.4 Vision and text embedding space recalibration
Although OpenAI CLIP and its derivative works [38, 50] achieve high accuracy in matching text with image patch, they focus on noun (such as “wine glass” and “cup”) instead of visual attributes (such as “transparency” and “narrow rim”). To match the image patches with task-specific visual attribute words, it is pivotal to re-calibrate the embedding space of vision and text. Therefore we design a transformer-based aligner module, next to the pre-trained VLM. In Fig. 3(a), we pass the vision embedding matrix and text embedding matrix into a multi-layer aligner module to generate the new re-calibrated vision and text embedding matrix, $\textbf{E}^{\textbf{m}}_{\textbf{V}}$ and $\textbf{E}^{\textbf{m}}_{\textbf{T}}$ respectively. Here m represents the number of layers in aligner module. In Fig. 3(b), we present the network architecture of each layer of the aligner module. There are multiple layers of aligner module where each layer has two self-attention modules, one cross-attention module, and one feed-forward network layer. For the $\textbf{i}^{\textbf{th}}$ aligner module layer, the input ($\textbf{E}^{\textbf{i}}_{\textbf{V}}$ and $\textbf{E}^{\textbf{i}}_{\textbf{T}}$) is the vision and text embedding matrix coming from last aligner module layer. At each layer, both vision and text embedding matrix will first go through a self-attention layer. The purpose of the self-attention layer is to learn the most optimal affordance that could be used to solve this task. For example, as is shown in Fig. 1(a) Scene 5, to solve the task of sit comfortably we prefer to choose chair over bed. After self-attention layer, we use a cross-attention layer to recalibrate the text embedding and vision embedding. The last part of the aligner network is a fully connected feed forward layer where the text embedding matrix will pass through. Like previous works that try to fine-tune pre-trained vision-language model [10, 15], in our aligner module, we focus on the text embedding vectors transformation. Therefore, we only pass text embedding matrix through the last feed forward layer.